OpenStudy (momo):

A point Z on is chosen at random, and W is between X and Y. If XY = 8 and XW = 2, what is the probability that Z is on ?

OpenStudy (anonymous):

What is Z on?

OpenStudy (momo):

oh XW

OpenStudy (anonymous):

8/100 = 2/x x = 25% would that be the answer?

OpenStudy (momo):

that is one of my options if you make it a decimal. .25